### v1

@@ 5.1 Methodology
We mainly consider two baselines: ZoneWrite-Only and ZoneAppend-Only. ZoneWrite-Only always uses Zone Write for all writes (i.e., $G=1$), while ZoneAppend-Only always uses Zone Append for all writes and does not employ a group-based data layout (i.e., $G=S$) (§3.2). We also compare ZapRAID with a simplified version of RAIZN [42], which we implement as a user-space block device based on SPDK (referred to as RAIZN-SPDK) for single-segment writes. Note that the open-source version of RAIZN [5] is currently implemented as a kernel device mapper, so we re-implement RAIZN as a user-space block device for fair comparisons with ZapRAID.
@@ 5.2 Microbenchmarks
We examine the performance of ZapRAID in writes, normal reads, degraded reads, crash recovery, and full-drive recovery. In our default setup, we configure the logical space capacity as 200 GiB and set $G=256$ for ZapRAID in its group-based data layout. We focus on one open segment with various chunk sizes. We do not consider garbage collection (by limiting the total write size to be less than the logical space capacity) and keep the L2P table entirely in memory. Note that we also consider various group sizes (Exp#3), crash recovery and full-drive recovery performance with different logical space capacities (Exp#5), multiple open segments (Exp#7), the overhead of garbage collection (Exp#8), and the overhead of L2P table offloading (Exp#9).
@@
Compared with RAIZN-SPDK, all other three schemes show much higher performance on a single segment. The reasons are two-fold. First, RAIZN does not support Zone Append for intra-zone parallelism. Second, it uses partial parity updates to acknowledge individual write requests instead of issuing full-stripe writes as in the other three schemes. This limits the concurrency of write requests and hence the overall write performance.
@@
Our current testbed has limited parallelism as it only has four real ZNS SSDs. To examine the scalability of ZapRAID with a higher degree of parallelism, we have also built an emulated ZNS SSD array with FEMU [49]. We set up a (6+1)-RAID-5 array on seven FEMU SSDs, using 4$\times$128-GiB Intel Optane Persistent Memory Modules [2] in Memory Mode as the emulated persistent storage backend. Each FEMU SSD is configured with four channels with four chips each. Each chip has 64 flash blocks with 4,096 4-KiB flash pages each. Each zone spans 16 chips, meaning that I/Os can be issued to up to 16 chips in parallel. We configure the page write latency as 140 $\mu$s [50] and do not emulate write buffers in SSDs that can reduce write latency. The page write latency translates to the write throughput of 27.9 MiB/s per chip, or equivalently, 446.4 MiB/s per zone if the parallelism is fully exploited. Since the FEMU SSD does not support flash-level page metadata, we store the page metadata in an in-memory array. We configure seven FEMU SSDs in a FEMU virtual machine, installed with Ubuntu 22.04, 32-GiB memory, and 12 vCPUs. We export the ZapRAID prototype as a block device with 64 GiB of logical space. We use FIO to write 64 GiB of data with a block size of 4 KiB to ZapRAID in the virtual machine for various chunk sizes and queue depths.
@@
 Refer to caption Refer to caption
 (a) Real SSDs (b) FEMU SSDs
Figure 11: Exp#6 (Scalability).
@@
We evaluate the write performance ZapRAID on multiple open segments based on hybrid data management (§3.3) by varying the numbers of small-chunk and large-chunk segments (i.e., $N_{s}$ and $N_{l}$, respectively). We fix $C_{s}=$ 8 KiB, $C_{l}=$ 16 KiB, and the number of open segments as four. We consider different combinations of $N_{s}$ and $N_{l}$, such that $N_{s}+N_{l}=4$. Note that if $N_{s}=0$ (resp. $N_{l}=0$), the small chunks (resp. the large chunks) are issued to large-chunk (resp. small-chunk) segments.
@@
Figures 12(a)-12(d) show the throughput results. ZoneWrite-Only and ZoneAppend-Only cannot always achieve the highest throughput in all cases. For example, ZoneAppend-Only has 65.7% higher throughput than ZoneWrite-Only for 4-KiB writes and $N_{s}=1$ due to better intra-zone parallelism under small writes (Figure 12(a)), but ZoneWrite-Only has 27.2% higher throughput than ZoneAppend-Only for 16-KiB writes and $N_{s}=1$ (Figure 12(c)). On the other hand, ZapRAID achieves the highest throughput, or similar throughput to the best scheme, in all cases by carefully combining Zone Append and Zone Write.
@@
Figure 12: Exp#7 (Impact of $N_{s}$ and $N_{l}$). We show the throughput (figures (a)-(d)) and 95th-percentile latencies (figures (e)-(h)) under four open segments.
@@
Figure 13 shows the results. As expected, the write throughput increases (i.e., the garbage collection overhead is less) with more reserved space and with more skewed or sequential workloads. Since real-world workloads are often skewed in practice (e.g., block storage [70] and key-value [16, 22] workloads), we expect that the garbage collection overhead is acceptable in ZapRAID. Note that the write throughput under multiple open segments (Figure 13(c)) is higher than under single open segments (Figures 13(a) and 13(b)), meaning that the garbage collection overhead can be reduced via inter-zone parallelism.
@@
We evaluate the overhead of offloading the L2P table entries to ZNS SSDs (§3.1). We consider the same four write workloads as in Exp#8 and focus on the setting of $(N_{s},N_{l})=(2,2)$ with four open segments. We set the physical space capacity as 300 GiB and issue writes to 200 GiB of data (i.e., the reserved space is 50%). We vary the maximum memory size allocated for the in-memory L2P table from 50 MiB to 200 MiB. Recall that each L2P table entry has a size of four bytes for each 4-KiB block (§3.1). Thus, the in-memory L2P table can store the entries from 50 GiB to 200 GiB of data. Note that when the in-memory L2P table is allocated with 200 MiB, it can keep all L2P table entries in memory (i.e., no offloading).
@@ 5.3 Trace-Driven Experiments
Figures 15(a) and 15(b) show the throughput results. For a single open segment with 4-KiB and 16-KiB chunks (Figure 15(a)), the throughput of ZapRAID is 69.4% and 6.4% higher than that of ZoneWrite-Only, respectively, and is close to that of ZoneAppend-Only; the findings are consistent with those in Exp#1. For multiple open segments, the findings are also consistent with those in Figure 12(d) of Exp#7. For $(N_{s},N_{l})=(2,2)$, all three schemes have similar throughput, while for $(N_{s},N_{l})=(1,3)$, ZapRAID has 25.3% higher throughput than ZoneWrite-Only and similar throughput to ZoneAppend-Only.
@@
  (a) Throughput (single segment) (b) Throughput (multiple segments) (c) 95p latency (single segment) (d) 95p latency (hybrid segments)
Figure 15: Exp#10 (Cloud block storage workloads).
Figures 15(c) and 15(d) show the 95th-percentile latencies, in which the trends are also consistent with Exp#1 and Exp#7, respectively. For a single open segment with 4-KiB and 16-KiB chunks (Figure 15(c)), ZapRAID has 56.6% lower and 39.6% higher 95th-percentile latencies than ZoneWrite-Only under the chunk sizes of 4 KiB and 16 KiB, respectively, and comparable 95th-percentile latencies with ZoneAppend-Only. For multiple open segments (Figure 15(d)), all three schemes have similar latencies for $(N_{s},N_{l})=(2,2)$, while ZoneWrite-Only has a higher 95th-percentile latency for $(N_{s},N_{l})=(1,3)$ as its performance is bottlenecked by the small writes that are forwarded to the only small-chunk open segment.

### v2

@@ 5.1 Methodology
We mainly consider two baselines: ZoneWrite-Only and ZoneAppend-Only. ZoneWrite-Only always uses Zone Write for all writes (i.e., $G=1$), while ZoneAppend-Only always uses Zone Append for all writes and does not employ a group-based data layout (i.e., $G=S$) (§3.2). For fair comparisons, when there are multiple open segments, ZoneWrite-Only and ZoneAppend-Only also write small chunks to small-chunk segments and large chunks to large-chunk segments, only using Zone Write and Zone Append, respectively.
We also compare ZapRAID with a simplified version of RAIZN [45], which we implement as a user-space block device based on SPDK (referred to as RAIZN-SPDK). Note that the open-source version of RAIZN [6] is currently implemented as a kernel device mapper and exports a ZNS interface as a logical ZNS SSD volume, so we re-implement RAIZN as a user-space block device for fair comparisons with ZapRAID. Specifically, the original RAIZN reserves at least three additional metadata zones (in addition to regular zones for data storage) in each drive, including one zone for appending partial parity updates before acknowledging each write request, one zone for appending metadata to ensure write atomicity, and at least one zone for the garbage collection of metadata zones. It caches the written data in an in-memory stripe buffer and writes the partial parity updates to the partial-parity-dedicated metadata zone, so as to support crash recovery even if a system crash happens before the whole stripe is written. When the stripe buffer is full, RAIZN generates and writes the complete parity chunks to the regular zones, and discards the corresponding partial parity updates. RAIZN differs from ZoneWrite-Only, ZoneAppend-Only, and ZapRAID in that it acknowledges each write request after the corresponding partial parity updates are appended to the dedicated metadata zone, while the other three schemes leverage concurrent write requests to form full stripes before acknowledging the write requests.
In our RAIZN-SPDK implementation, we focus on achieving high performance and simplify our implementation by removing the metadata zones reserved for write atomicity and garbage collection. RAIZN-SPDK also reserves two metadata zones for high performance. Specifically, it writes partial parity updates to one of the two metadata zones. When the metadata zone becomes full, RAIZN-SPDK resets the zone and writes the partial parity updates to another metadata zone in parallel. For high performance, RAIZN-SPDK uses Zone Append to write partial parity updates to a metadata zone as described in [45]. When there are multiple open segments, RAIZN-SPDK uses Zone Write to write small chunks to small-chunk segments and large chunks to large-chunk segments as in ZoneWrite-Only, while using Zone Append to write partial parity updates of all segments to a metadata zone.
We also attempt to fairly compare ZapRAID with the conventional RAID array that runs on traditional block-interface SSDs. However, the firmware on our ZN540 SSDs is locked and cannot be directly modified, so we cannot make comparisons on the same hardware as in [45]. Thus, we leverage FEMU-emulated SSDs [53] for comparisons and build the conventional RAID array with mdadm [16] (see Exp#6 for details).
Our experiments consider the configurations with a single open segment and multiple open segments. If there is a single open segment, ZapRAID always issues writes via Zone Append and applies the group-based data layout; if there are multiple open segments, ZapRAID enables hybrid segment management and issues writes via a combination of Zone Append and Zone Write (the exact configurations are further explained in individual experiments).
@@ 5.2 Microbenchmarks
We examine the performance of ZapRAID in writes, normal reads, degraded reads, crash recovery, and full-drive recovery. In our default setup, we configure the logical space capacity as 200 GiB and set $G=256$ for ZapRAID in its group-based data layout. First, we focus on one open segment with various chunk sizes in Exp#1-Exp#6. We do not consider garbage collection (by limiting the total write size to be less than the logical space capacity) and keep the L2P table entirely in memory. We also consider the impact of various group sizes (Exp#3), the impact of RAID schemes (Exp#4), as well as the crash recovery and full-drive recovery performance with different logical space capacities (Exp#5). Furthermore, we consider multiple open segments in Exp#7-Exp#9, including the write performance under multiple open segments (Exp#7), the overhead of garbage collection (Exp#8), and the overhead of L2P table offloading (Exp#9).
@@
Compared with RAIZN-SPDK, all other three schemes show much higher performance on a single segment. The reasons are two-fold. First, RAIZN does not support Zone Append for intra-zone parallelism. Second, it uses partial parity updates to acknowledge individual write requests instead of issuing full-stripe writes as in the other three schemes. This limits the concurrency of write requests and hence the overall write performance. In Exp#7, we further compare ZapRAID with RAIZN-SPDK under multiple open segments.
@@
Our current testbed has limited parallelism as it only has four real ZNS SSDs. To examine the scalability of ZapRAID with a higher degree of parallelism, we have also built an emulated ZNS SSD array with FEMU [53]. We set up a (6+1)-RAID-5 array on seven FEMU SSDs, using 4$\times$128-GiB Intel Optane Persistent Memory Modules [2] in Memory Mode as the emulated persistent storage backend. Each FEMU SSD is configured with four channels with four chips each. Each chip has 64 flash blocks with 4,096 4-KiB flash pages each. Each zone spans 16 chips, meaning that I/Os can be issued to up to 16 chips in parallel. We configure the page write latency as 140 $\mu$s [54] and do not emulate write buffers in SSDs that can reduce write latency. The page write latency translates to the write throughput of 27.9 MiB/s per chip, or equivalently, 446.4 MiB/s per zone if the parallelism is fully exploited. Since the FEMU SSD does not support flash-level page metadata, we store the page metadata in an in-memory array. We configure seven FEMU SSDs in a FEMU virtual machine, installed with Ubuntu 22.04, 32-GiB memory, and 12 vCPUs. We export the ZapRAID prototype as a block device with 96 GiB of logical space. We use FIO to issue the random writes of 64 GiB of data with a block size of 4 KiB to ZapRAID in the virtual machine for various chunk sizes and queue depths.
@@
We also examine the performance of the conventional RAID array based on FEMU-emulated block-interface SSDs. Similar to FEMU ZNS SSDs, we configure a (6+1)-RAID-5 array on seven FEMU black-box SSDs with Linux mdadm [16]. The configurations of flash blocks, channels, and chips are the same as FEMU ZNS SSDs, and each FEMU SSD has 16 GiB of physical space. We also keep the same page write latency and the same settings as the FEMU virtual machine. We export the RAID array as a block device with 64 GiB of logical space, so the overprovisioned space is about $(\frac{16\times 6}{64}-1)\times 100\%=50\%$. We use FIO to issue random writes of 64 GiB to the RAID array for various chunk sizes and queue depths. Figure 11(c) shows the results. Compared with ZapRAID in Figure 11(b), the throughput of the conventional RAID array is much lower (up to 200 MiB/s only). The results are expected, since updating a data chunk in the conventional RAID array issues a read-modify-write operation, which reads the current data chunk being modified and the current parity chunk, modifies the parity chunk, and writes back the new data chunk and new parity chunk. In contrast, ZapRAID issues concurrent write requests to form full stripes and avoids the overhead of parity updates.
Refer to caption Refer to caption Refer to caption
(a) ZapRAID on real ZNS SSDs (b) ZapRAID on FEMU ZNS SSDs (c) Conventional RAID on FEMU block-interface SSDs
Figure 11: Exp#6 (Scalability). The results in figures (a) and (b) are based on ZapRAID, and the results in figure (c) are based on the conventional RAID array.
@@
We evaluate the write performance ZapRAID on multiple open segments based on hybrid data management (§3.3) by varying the numbers of small-chunk and large-chunk segments (i.e., $N_{s}$ and $N_{l}$, respectively). We fix $C_{s}=$ 8 KiB, $C_{l}=$ 16 KiB, and the number of open segments as four. We consider different combinations of $N_{s}$ and $N_{l}$, such that $N_{s}+N_{l}=4$. Note that if $N_{s}=0$ (resp. $N_{l}=0$), all writes are issued to large-chunk (resp. small-chunk) segments.
@@
Figures 12(a)-12(d) show the throughput results. ZoneWrite-Only and ZoneAppend-Only cannot always achieve the highest throughput in all cases. For example, ZoneAppend-Only has 65.7% higher throughput than ZoneWrite-Only for 4-KiB writes and $(N_{s},N_{l})=(1,3)$ due to better intra-zone parallelism under small writes (Figure 12(a)), but ZoneWrite-Only has 27.2% higher throughput than ZoneAppend-Only for 16-KiB writes and $(N_{s},N_{l})=(1,3)$ (Figure 12(c)). On the other hand, ZapRAID achieves the highest throughput, or similar throughput to the best scheme, in all cases by carefully combining Zone Append and Zone Write.
@@
Figure 12: Exp#7 (Multiple open segments). We show the throughput (figures (a)-(d)) and 95th-percentile latencies (figures (e)-(h)) under four open segments.
@@
[…]
Refer to caption Refer to caption
(a) $N_{l}=2$ and varying $N_{s}$, throughput (b) $N_{s}=2$ and varying $N_{l}$, throughput
Refer to caption Refer to caption
(c) $N_{l}=2$ and varying $N_{s}$, 95p latency (d) $N_{s}=2$ and varying $N_{l}$, 95p latency
Figure 13: Exp#7 (Multiple open segments). We compare the throughput and 95th-percentile latencies of RAIZN-SPDK and ZapRAID under hybrid 75% 4-KiB and 25% 16-KiB writes and different combinations of $(N_{s},N_{l})$.
We also compare ZapRAID with RAIZN-SPDK under multiple open segments. We focus on the workloads with hybrid 75% 4-KiB and 25% 16-KiB writes, and examine different combinations of $(N_{s},N_{l})$. We consider two settings: (i) we fix $N_{l}=2$ and vary $N_{s}$ from 0 to 6, and (ii) we fix $N_{s}=2$ and vary $N_{l}$ from 0 to 6.
Figure 13(a) shows the throughput results of RAIZN-SPDK and ZapRAID when we fix $N_{l}=2$ and vary $N_{s}$. For RAIZN-SPDK, when $(N_{s},N_{l})$ changes from $(0,2)$ to $(1,2)$, its throughput drops from 440.4 MiB/s to 268.6 MiB/s. The reason is that RAIZN-SPDK forwards all 4-KiB writes to two large-chunk segments for $(N_{s},N_{l})=(0,2)$, but for $(N_{s},N_{l})=(1,2)$, it forwards all 4-KiB writes to only one small-chunk segment and fails to leverage inter-zone parallelism for these small writes. When $(N_{s},N_{l})$ changes from $(1,2)$ to $(6,2)$, the throughput of RAIZN-SPDK increases to 1,077.6 MiB/s with the increasing inter-zone parallelism across small-chunk open segments. For ZapRAID, with the help of intra-zone parallelism for small writes, it maintains similar throughput for $(N_{s},N_{l})=(0,2)$ and $(N_{s},N_{l})=(1,2)$, even though the latter has only one small-chunk open segment.
Figure 13(b) shows the throughput results of RAIZN-SPDK and ZapRAID when we fix $N_{s}=2$ and vary $N_{l}$. The throughput of RAIZN-SPDK is 444.6 MiB/s for $(N_{s},N_{l})=(2,0)$, but only increases to 488.4-498.0 MiB/s when $(N_{s},N_{l})$ varies from $(2,1)$ to $(2,6)$. The reason is that most of the write requests (75% in total) are 4-KiB writes, and the performance is bottlenecked by the limited number of small-chunk open segments.
Figures 13(c) and 13(d) show the 95th-percentile latency results of RAIZN-SPDK and ZapRAID, and the observations are consistent with those in the throughput results. Overall, when there are multiple open segments, ZapRAID still achieves significantly higher throughput and lower latencies than RAIZN-SPDK, which incurs large overhead of partial parity updates.
($N_{s},N_{l}$) (0, 2) (1,2) (2,2) (6,2) (2,0) (2,1) (2,2) (2,6)
ZapRAID Wait phase 27.01 31.92 34.38 41.01 39.10 35.89 34.38 38.17
Data phase 72.86 74.30 66.86 67.25 76.31 66.23 66.86 70.27
Parity phase 5.44 3.69 3.68 3.52 3.43 3.76 3.68 3.60
RAIZN-SPDK Wait phase 775.46 1282.24 678.87 292.05 861.28 694.47 678.87 682.73
Data phase 23.66 22.62 23.96 31.46 26.08 23.83 23.96 24.24
Parity phase 1.01 1.02 1.02 1.11 1.04 1.04 1.02 1.04
Table 1: Exp#7 (Multiple open segments). Breakdown of latencies (in microseconds) of ZapRAID and RAIZN-SPDK for Figure 13.
Table 1 further shows the latency breakdown of RAIZN-SPDK and ZapRAID based on some of the settings of $(N_{s},N_{l})$ in Figure 13, so as to explain the large overhead of partial parity updates in RAIZN-SPDK. For each write request, we record four timestamps: (i) $t_{1}$, the time when the request is issued by fio, (ii) $t_{2}$, the time when the request starts writing data chunks, (iii) $t_{3}$, the time when the request finishes writing data chunks, and (iv) $t_{4}$, the time when the request completes writing all parity chunks (including partial parity updates for RAIZN-SPDK). We divide each write request into three phases: (i) the wait phase (i.e., $t_{2}-t_{1}$), (ii) the data phase (i.e., $t_{3}-t_{2}$), and (iii) the parity phase (i.e., $t_{4}-t_{3}$). The table shows that RAIZN-SPDK has a prolonged wait phase, as each of its write requests must update the parity chunks and wait for the completion of the partial parity updates of the previous write request for the same segment, thereby limiting the concurrency of writes. In contrast, ZapRAID has a short wait phase as it can issue writes concurrently (recall that our default queue depth is 64). Note that ZapRAID has longer data and parity phases than RAIZN-SPDK since its concurrent writes may overlap, so on average each write request takes longer to complete. Nevertheless, the overall write throughput of ZapRAID is significantly higher than that of RAIZN-SPDK.
@@
Figure 14 shows the results. As expected, the write throughput increases (i.e., the garbage collection overhead is less) with more reserved space and with more skewed or sequential workloads. Since real-world workloads are often skewed in practice (e.g., block storage [74] and key-value [18, 24] workloads), we expect that the garbage collection overhead is acceptable in ZapRAID. Note that the write throughput under multiple open segments (Figure 14(c)) is higher than under single open segments (Figures 14(a) and 14(b)), meaning that the garbage collection overhead can be reduced via inter-zone parallelism.
@@
We evaluate the overhead of offloading the L2P table entries to ZNS SSDs (§3.1). We consider the same four write workloads as in Exp#8 and focus on the setting of $(N_{s},N_{l})=(2,2)$ with four open segments. We set the physical space capacity as 300 GiB and issue writes to 200 GiB of data (i.e., the overprovisioned space is 50%). We vary the maximum memory size allocated for the in-memory L2P table from 50 MiB to 200 MiB. Recall that each L2P table entry has a size of four bytes for each 4-KiB block (§3.1). Thus, the in-memory L2P table can store the entries from 50 GiB to 200 GiB of data. Note that when the in-memory L2P table is allocated with 200 MiB, it can keep all L2P table entries in memory (i.e., no offloading).
@@ 5.3 Trace-Driven Experiments
Figures 16(a) and 16(b) show the throughput results. For a single open segment with 4-KiB and 16-KiB chunks (Figure 16(a)), the throughput of ZapRAID is 69.4% and 6.4% higher than that of ZoneWrite-Only, respectively, and is close to that of ZoneAppend-Only; the findings are consistent with those in Exp#1. For multiple open segments, the findings are also consistent with those in Figure 12(d) of Exp#7. For $(N_{s},N_{l})=(2,2)$, all three schemes have similar throughput, while for $(N_{s},N_{l})=(1,3)$, ZapRAID has 25.3% higher throughput than ZoneWrite-Only and similar throughput to ZoneAppend-Only.
@@
(a) Throughput (single segment) (b) Throughput (multiple segments) (c) 95p latency (single segment) (d) 95p latency (multiple segments)
Figure 16: Exp#10 (Cloud block storage workloads). Average throughput and 95 latency results of all 20 volumes in Alibaba Cloud traces.
Figures 16(c) and 16(d) show the 95th-percentile latencies, in which the trends are also consistent with Exp#1 and Exp#7, respectively. For a single open segment with 4-KiB and 16-KiB chunks (Figure 16(c)), ZapRAID has 56.6% lower and 39.6% higher 95th-percentile latencies than ZoneWrite-Only under the chunk sizes of 4 KiB and 16 KiB, respectively, and comparable 95th-percentile latencies with ZoneAppend-Only. For multiple open segments (Figure 16(d)), all three schemes have similar latencies for $(N_{s},N_{l})=(2,2)$, while ZoneWrite-Only has a higher 95th-percentile latency for $(N_{s},N_{l})=(1,3)$ as its performance is bottlenecked by the small writes that are forwarded to the only small-chunk open segment.
Volume No. 1 2 3 4 5 6 7 8 9 10
Ratios (%) Small writes 83.35 83.06 83.30 83.56 83.29 83.50 83.03 83.36 82.66 82.98
Large writes 1.63 1.64 1.67 1.70 1.75 1.76 1.77 1.78 1.83 1.86
Thpt (MiB/s) (4 KiB chunks) ZoneWrite-Only 1,023.7 947.0 1,035.1 1,001.5 937.5 981.5 1,035.3 968.2 986.2 988.7
ZapRAID 1,740.6 1,711.3 1,702.4 1,699.1 1,660.1 1,657.9 1,612.7 1,619.4 1,681.0 1,636.3
Thpt (MiB/s) (($N_{s},N_{l}$)=(1,3)) ZoneWrite-Only 2,486.4 2,341.2 2,323.1 2,401.3 2,403.7 2,333.7 2,311.0 2,360.3 2,344.8 2,410.5
ZapRAID 3,175.9 3,225.1 3,271.9 3,193.7 3,183.1 3,227.5 3,231.6 3,205.8 3,249.7 3,298.2
Volume No. 11 12 13 14 15 16 17 18 19 20
Ratios (%) Small writes 82.96 86.68 84.00 81.58 80.98 81.41 72.45 73.85 65.78 63.05
Large writes 1.92 1.95 3.42 3.70 4.49 10.33 16.83 20.95 21.11 24.89
Thpt (MiB/s) (4 KiB chunks) ZoneWrite-Only 948.4 1,007.2 982.2 995.6 989.2 972.1 970.4 993.0 1,004.0 971.5
ZapRAID 1,627.3 1,710.9 1,634.0 1,645.8 1,699.1 1,666.3 1,684.8 1,736.4 1,656.4 1,665.6
Thpt (MiB/s) (($N_{s},N_{l}$)=(1,3)) ZoneWrite-Only 2,372.7 2,399.1 2,412.2 2,439.7 2,593.6 2,637.1 2,703.2 2,773.4 2,862.1 2,825.4
ZapRAID 3,175.9 3,225.1 3,271.9 3,193.7 3,183.1 3,227.5 3,231.6 3,205.8 3,249.7 3,298.2
Table 2: Exp#10 (Cloud block storage workloads). Statistics of the 20 individual volumes in Alibaba Cloud traces.
We also show the statistics of individual volumes by presenting the ratios of numbers of small writes (i.e., 4 KiB) and large writes (i.e., at least 16 KiB) over the total number of write requests as well as the throughput of each volume. For brevity, since all three schemes have close throughput under one single open segment with 16-KiB chunks and four open segments with $(N_{s},N_{l})=(2,2)$, we only present the results of one single open segment with 4-KiB chunks and four open segments with $(N_{s},N_{l})=(1,3)$. Also, we only focus on the results under ZoneWrite-Only and ZapRAID to show their variations across the volumes.
Table 2 shows the results. We sort the volumes by the ratio of large write requests. For one single open segments under 4-KiB chunks, the throughput is close over all volumes for each scheme. For four open segments with $(N_{s},N_{l})=(1,3)$, ZoneWrite-Only shows lower throughput for a lower ratio of large writes; for example, the throughput results of Volumes 2 and 20 are 2,341.2 MiB/s and 2,825.4 MiB/s with large-write ratios of 1.64% and 24.89%, respectively. The reason is that with fewer large writes, more write requests are written to the only $N_{s}=1$ small-chunk open segment, and ZoneWrite-Only fails to utilize intra-zone parallelism to issue writes efficiently. In contrast, ZapRAID maintains higher throughput with an increase of 14.7-40.8% compared with ZoneWrite-Only under $(N_{s},N_{l})=(1,3)$, across all 20 volumes.